OpenStudy (anonymous):

Confused can someone help me thanks. Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take Carol to catch up to Harold

OpenStudy (jamesj):

Harolds position at time time t, H(t) = 55t, make sense? And Carol's position at time t (for t > 2 hours) is C(t) = 65(t-2). So far so good?

OpenStudy (anonymous):

yea thanks so much

OpenStudy (jamesj):

Now you want to know when Harold's position is equal to Carol's position. That is when Carol catches up. I.e., you want to know the value of t such that H(t) = C(t)

OpenStudy (anonymous):

I am confused now sorry

OpenStudy (jamesj):

How far has Harold travelled after t hours?

OpenStudy (anonymous):

55m

OpenStudy (jamesj):

He's traveling at 55 miles/hour. Hence after t hours, he has travelled a distance of (55 miles/hour) x (t hours) = 55t miles.

OpenStudy (jamesj):

Now Carol. At time t = 2 hours, she leaves. At t = 2 hours exactly, she hasn't travelled anywhere at all. How far has she travelled at t = 3 hours?

OpenStudy (anonymous):

5 hours after? sorry

OpenStudy (jamesj):

Try and follow the logic I'm building up for you. Carol, when she is traveling, travels at a speed of 65 miles/hour. Hence at t = 3 hours, she has been traveling for one hour. How far has she traveled at t = 3 hours?

OpenStudy (anonymous):

21

OpenStudy (jamesj):

At t = 3 hours, she's been driving for one hour at 65 mph. Hence she has travelled a distance of distance = speed x time = 65 miles/hour x 1 hour = 65 miles Make sense?

OpenStudy (anonymous):

I got 65/2*1 =32.5 did I do it right I am very

OpenStudy (anonymous):

sorry to bother u

OpenStudy (jamesj):

Let me as you, at time t, for t > 2 hours, how long has Carol been driving?

OpenStudy (jamesj):

Do you agree she's been driving (t-2) hours?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Hence at time t, for t > 2, how far has Carol travelled?

OpenStudy (anonymous):

=32m 65/2

OpenStudy (jamesj):

No. Distance = Speed x Time Her speed is 65 miles/hour. And the time she's been driving is (t-2) hours. Hence how far has she driven at time t?

OpenStudy (anonymous):

130 miles

OpenStudy (jamesj):

No. Where did t go?

OpenStudy (anonymous):

I am confusing my self with t and miles

OpenStudy (jamesj):

Distance = Speed x Time = 65 miles/hour x (t-2) hours = 65(t-2) miles

OpenStudy (anonymous):

I am so sorry

OpenStudy (anonymous):

thanks for all your help I am so sorry to bother u

OpenStudy (jamesj):

We're going to finish this. Does that formula for Carol's distance make sense to you? Distance = Speed x Time = 65 miles/hour x (t-2) hours = 65(t-2) miles

OpenStudy (anonymous):

no sorry

OpenStudy (jagatuba):

JamesJ, would it be easier to simplify the problem in to simpler terms that she may understand better?

OpenStudy (jamesj):

At time t = 3, how long has Carol been driving?

OpenStudy (jagatuba):

She seems to be having a hard time grasping the concept of distance traveled over time.

OpenStudy (jamesj):

( I don't know how to get away from that. That's the basic concept we're measuring here. )

OpenStudy (jamesj):

At time t = 3, how long has Carol been driving?

OpenStudy (jagatuba):

Or he (sorry I made an assumption with no basis. No offense).

OpenStudy (anonymous):

I am a she thanks for helping

OpenStudy (anonymous):

I am very sorry to bother I just dont really understand this problem

OpenStudy (jagatuba):

It's okay. That's what this website is for

OpenStudy (jagatuba):

Let's try a different approach.

OpenStudy (anonymous):

okay

OpenStudy (jagatuba):

If you are traveling at 65 miles per hour, how far will you have gone in 1 hour?

OpenStudy (anonymous):

do I multiply 65 * 60 ?

OpenStudy (jagatuba):

no

OpenStudy (anonymous):

65 miles per hour means you travel 65 miles in 1 (one) hour. In two hours you travel 2*65 miles and so on.

OpenStudy (jagatuba):

65 miles per hour can also be thought of as 65 miles per 1 hour. So in one hour you will have traveled 65 miles. Does that make sense?

OpenStudy (jagatuba):

Right Prebz

OpenStudy (anonymous):

130 ?

OpenStudy (jagatuba):

At 65 mph you will have traveled 130 miles in 2 hours does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (jagatuba):

Okay so we know that Harold left two hours before Carol right? He has a 2 hour head start in Carol.

OpenStudy (anonymous):

okay

OpenStudy (jagatuba):

He is going 55 mph so how far has he already traveled when Carol finally gets going?

OpenStudy (anonymous):

55*2 will equal 110

OpenStudy (jagatuba):

Correct.

OpenStudy (jagatuba):

Now at any point during their travels the time difference between Carol and Harold is going to be the same. Do you agree?

OpenStudy (anonymous):

no theres will be different?

OpenStudy (jagatuba):

How so?

OpenStudy (jagatuba):

There is no way that Carol can change the time difference without a time machine. She will always into infinity haver left 2 hours after Harold.

OpenStudy (anonymous):

yes sorry I was thinking in a different way sorry ur right

OpenStudy (jagatuba):

Ok so let's call this time difference t-2

OpenStudy (jagatuba):

So let's look at what that means real quick so, I'm sure you understand.

OpenStudy (anonymous):

k

OpenStudy (jagatuba):

t = the time Harold has driven and t-2 = the time Carol has driven. So when harord has driven for 3 hours how long has Carol driven?

OpenStudy (anonymous):

65*3 195?

OpenStudy (anonymous):

Harold would have driven 55*3 miles, Carol on the other hand would have driven 65(3-2)miles

OpenStudy (jagatuba):

no. at this point we are just looking at time only. We will get to distance in a minute. Since time for Harold is t and time for Carol is t-2 then to figure out how long Carol has driven all you have to do is substitute in how long Harold has driven. In this case he has driven 3 hours so Carol has driven 3-2 hours. Does the make sense?

OpenStudy (jagatuba):

Yes Prebz

OpenStudy (anonymous):

yea

OpenStudy (jagatuba):

Okay cool. So now you understand that time is a constant in this equation and that constant is t-2. Right?

OpenStudy (anonymous):

yes

OpenStudy (jagatuba):

Okay now we can deal with speed and distance which I think you are finally getting the idea of. You know that by the time Carol starts Harold has driven 55*2 miles right?

OpenStudy (jagatuba):

That is 55 mph * 2 hours.

OpenStudy (anonymous):

yes and harold is 110 55*2

OpenStudy (jagatuba):

Right. Now what we want to know is how long it takes Carol to catch up to Harold. So we can express this as an equation. 55(t)=65(t-2) Does that makes sense to you?

OpenStudy (anonymous):

yes it does

OpenStudy (jamesj):

Does it? Where does that equation come from?

OpenStudy (jagatuba):

Mmm. I think so. Am I mistaken?

OpenStudy (jamesj):

It's definitely right. I'm just asking if lala understands where it comes from.

OpenStudy (jagatuba):

Ah right.

OpenStudy (jagatuba):

BRB I have to go get my son from school.

OpenStudy (anonymous):

okay let me see If I got this right 55(110)=65(130-2) if not I give up so sorry to bother and take up your time thank you for ur patience in teaching me

OpenStudy (anonymous):

okay

OpenStudy (jagatuba):

Ok I'm back.

OpenStudy (jagatuba):

So my desire to not further confuse you caused me to rush to the equation 55t=65(t-2), but JamesJ was quick to point out that I did not explain how we got there. I don't want to further confuse you, so I'm going to break it down as simply as I know how and if you have further questions you can ask for clarification. We know that distance traveled (d) is equal to speed (s) multiplied by time (t) or d=st. Since we have two cars traveling, we have two equations: Harold: d=55t Carol: d=65(t-2) Since we are looking for the point where Carol catches up to Harold, we know that the distance is going to be equal in both equations. Therefore: 55t=65(t-2) And there you go. From here it is just a matter of solving for t, which gives you the amount of time in hours that it will take for Carol to catch up with Harold. It appears you have left the forum lalaland, so if you have any further questions or need more clarification just respond to this post and I or someone else can explain further until you understand. As I said before, don't worry about trouble us or putting us out. Everyone here is here to help each other, so asking and asking and asking until you understand is not going to put anyone out. As you can see from this thread, if one person is not getting through to you and explaining so that you understand, someone else can step in and try a different way of explaining things.

OpenStudy (anonymous):

thank you so much for helping me thank you sorry that I had to go but my internet quit on me. I went back and solved what you had taught me I understand know the problem. Thank you so much again.